I'm having problems with this can anyone help me?
Solve the system by graphing, if possible. (If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)
3x − 3y = 21
x − y = 13

Found 2 solutions by swincher4391, lynnlo:
Answer by swincher4391(1107)

Solve each in terms of y:
3x -3y = 21
3(x-y) = 21
x-y = 7
y = x-7
x-y = 13
-y = -x +13
y = x -13
When you graph these two lines you'll see they never intersect. Thus there is no solution and hence INCONSISTENT.

Answer by lynnlo(4176)

y=x-7
y=x-13
dependent,